# CQ-DINO: Mitigating Gradient Dilution via Category Queries for Vast Vocabulary Object Detection

### v3

@@ Abstract
With the exponential growth of data, traditional object detection methods are increasingly struggling to handle vast vocabulary object detection tasks effectively. We analyze two key limitations of classification-based detectors: positive gradient dilution, where rare positive categories receive insufficient learning signals, and hard negative gradient dilution, where discriminative gradients are overwhelmed by numerous easy negatives. To address these challenges, we propose CQ-DINO, a category query-based object detection framework that reformulates classification as a contrastive task between object queries and learnable category queries. Our method introduces image-guided query selection, which reduces the negative space by adaptively retrieving the top-K relevant categories per image via cross-attention, thereby rebalancing gradient distributions and facilitating implicit hard example mining. Furthermore, CQ-DINO flexibly integrates explicit hierarchical category relationships in structured datasets (e.g., V3Det) or learns implicit category correlations via self-attention in generic datasets (e.g., COCO). Experiments demonstrate that CQ-DINO achieves superior performance on the challenging V3Det benchmark (surpassing previous methods by 2.1% AP) while maintaining competitiveness on COCO. Our work provides a scalable solution for real-world detection systems requiring wide category coverage. The code is available at https://github.com/RedAIGC/CQ-DINO.
@@ 1 Introduction
With the rapid expansion of data, developing a robust AI system capable of large-scale object detection has become essential. This necessity is driven by the increasing complexity and diversity of real-world applications, where AI must manage an extensive vocabulary and dynamic environments. Vast vocabularies inherently present hierarchical category structures, as illustrated by classification datasets like ImageNet [7] and Bamboo [43]. Recent detection benchmarks, such as V3Det [35], featuring 13,204 object categories organized in hierarchical structures, highlight the magnitude of this challenge. While object detection has witnessed significant advancements [27, 17, 22, 10], scaling effectively to vast vocabularies remains a substantial challenge.
Category prediction mechanisms can be broadly categorized into four types, as illustrated in Fig. 1. Classification head-based methods employ feed-forward networks (FFNs) with sigmoid activation and Focal Loss [30] optimization. These approaches perform well on benchmarks with limited categories such as COCO [19] (80 categories) and Objects365 [31] (365 categories), but face fundamental challenges and scalability issues in vast vocabulary settings. Text-prompted contrastive methods leverage Vision-Language Models (VLMs) to encode target categories as text inputs, achieving strong open-vocabulary detection [40] performance. However, in vast vocabulary scenarios, text input sequences become prohibitively long, necessitating the splitting of category lists across multiple inference passes. This substantially limits their practical scalability and efficiency. Language model-generated methods approach open-ended [18] object detection by using language models to generate category labels without predefined candidate sets. Although these methods theoretically enable the detection of arbitrary object categories, they generally lack mechanisms to control the granularity of generated labels, which can result in substantial misalignment with practical detection requirements.
@@
To tackle these challenges, we introduce Category Query-based DINO (CQ-DINO), a novel architecture that encodes categories as learnable query embeddings. Our approach centers on image-guided query selection, which identifies relevant categories via category-to-image similarity through cross-attention. The key insight driving our method is that dynamic sparse category selection significantly reduces the negative search space. This selection mechanism provides three crucial benefits: (1) balancing the ratio between positive and negative gradients, (2) performing implicit hard mining by selecting the most similar categories, and (3) reducing memory and computational costs, making the framework scalable to extremely large vocabularies. Selected category queries interact with image features to generate object queries. From these object queries, bounding boxes are predicted through a cross-modality decoder. The final classifications are obtained using contrastive alignment between the object queries and the category queries. Unlike traditional classification head-based methods, our category query representation offers greater flexibility by naturally encoding inter-category relationships. For structured datasets with explicit category hierarchies like V3Det [35], we leverage the inherent tree structure to construct hierarchical category queries with an adaptive weighting mechanism that balances local and hierarchical features. For datasets without explicit hierarchies (e.g., COCO [19]), we employ self-attention mechanisms to implicitly learn the correlations between categories.
We evaluate CQ-DINO on both the vast vocabulary V3Det dataset [35] and the standard COCO benchmark [19]. Our approach surpasses previous state-of-the-art methods on V3Det while maintaining competitive results on COCO compared to DETR-based detectors. Our method benefits from vast vocabulary detection while maintaining competitive results in limited vocabulary scenarios.
@@ 2.1 Vast Vocabulary Object Detection
The progression of object detection benchmarks reflects a steady growth in category vocabulary, evolving from relatively small vocabulary datasets such as PASCAL VOC [9] (20 classes) and COCO [19] (80 classes), to larger vocabulary benchmarks, including Objects365 [31] (365 classes) and Open Images [15] (600 classes). Recently, Wang et al. [35] introduce V3Det, the first vast vocabulary object detection dataset, comprising 13,204 hierarchically structured categories. This unprecedented scale poses significant challenges in terms of scalability and representation.
Recent progress has been driven by the V3Det Challenge [34], yielding several methodological advancements. MixPLv2 proposes a semi-supervised framework that combines labeled V3Det data with unlabeled Objects365 [31] images through pseudo-labeling. RichSem-DINO-FocalNet enhances detection robustness by integrating the RichSem-DINO [24] framework with a FocalNet-Huge backbone [38] pretrained on Objects365. Most recently, Prova [4] introduces multi-modal image-text prototypes specifically optimized for V3Det’s fine-grained classification. However, these methods still rely on fundamentally similar classification architectures, raising questions about their effectiveness and scalability for even larger vocabularies beyond tens of thousands of categories.
@@ 2.2 Object Detectors
Classification Head-based Methods. Most object detection frameworks employ feedforward networks (FFNs) as fixed classification heads for category prediction. In two-stage detectors [27, 2, 11], region proposals are generated in a class-agnostic manner by Region Proposal Networks (RPNs), followed by FFN-based region classifiers. By contrast, one-stage detectors [30, 33, 8] predict bounding boxes and categories in a single step. Transformer-based methods, such as DETR [3], reformulate detection as a set prediction problem using learnable queries. Although DETR achieves an elegant end-to-end design, it suffers from slow convergence. Subsequent works [47, 36, 20, 16, 41, 25, 21, 12] mitigate these limitations. For example, Deformable DETR [47] proposes multi-scale deformable attention for sparse spatial sampling. DINO [41] improves performance via contrastive query denoising and mixed query selection. Despite these architectural advances, current methods fundamentally rely on classification heads with activation functions, typically optimized with Focal Loss [30] or Cross-Entropy Loss [44]. This design inherently constrains scalability and presents optimization challenges when extending to vast vocabulary detection scenarios.
@@
Text-prompted Contrastive Methods. Vision-language models have advanced the seamless integration of visual and textual modalities for open-vocabulary object detection. These methods encode target categories as text inputs and align visual and textual representations. For instance, GLIP [17] pioneers the use of contrastive learning between image regions and textual phrases. Grounding DINO [22] further improves cross-modal alignment through early fusion of vision and textual features. Similarly, DetCLIP [39] and RegionCLIP [45] leverage image-text pairs with pseudo-labels to enhance region-level semantic understanding and improve generalization. Despite these advances, text-prompted methods face scalability bottlenecks due to the limited capacity of text token inputs during inference. For example, GLIP [17] and Grounding DINO [22] restrict input prompts to approximately 128 tokens per pass, allowing for about 40 categories simultaneously. Thus, detecting all categories in a vast vocabulary benchmark like V3Det [35] (13,204 classes) would require over 331 sequential inference passes per image. This constraint renders current text-prompted methods computationally inefficient and impractical for real-time or large-scale detection.
@@
Language Model-Generated Methods. Advances in multimodal large language models (MLLMs) have inspired detection methods leveraging their visual understanding and generative abilities. Some MLLMs [5, 1, 32] show preliminary object detection abilities but exhibit limited localization precision and recognition granularity. To mitigate these limitations, recent works [18, 10, 14, 35, 37] use LLMs primarily as category generators rather than direct detectors. For instance, GenerateU [18] employs a T5-based decoder [6] to generate category names from visual features, reframing detection as text generation. Similarly, LLMDet [10] and ChatRex [14] utilize instruction-tuned LLMs to predict object categories from image features. While these generative approaches enable open-ended detection, they often produce inconsistencies due to limited controllability over label granularity. For example, given an image region of a “Persian cat”, the model may generate the generic term “cat”, causing semantic ambiguity and reduced accuracy for fine-grained detection tasks.
@@ 3.1 Challenges in Vast Vocabulary Detection
Existing methods for vast vocabulary object detection with $C$ categories ($C>10^{4}$), particularly those employing sigmoid-based classifiers with Focal Loss [30], face fundamental optimization challenges. We systematically analyze these issues through a simplified formulation using the Cross-Entropy Loss with sigmoid activation, revealing two critical limitations:
1) Positive Gradient Dilution. In vast vocabulary detection, the gradient signal for positive classes is overwhelmed by the aggregated negative gradients. Let $z_{c}$ denote the logit for class $c$ and $y_{c}\in\{0,1\}$ be its ground-truth label. The gradient of the Cross-Entropy Loss $\mathcal{L}$ with respect to $z_{c}$ is:
@@
$||\nabla_{z_{c^{+}}}\mathcal{L}||\ll\sum_{c^{-}\neq c^{+}}^{C}||\nabla_{z_{c^{-}}}\mathcal{L}||\approx(C-1)\cdot\epsilon,$ (2)
@@
$\rho=\frac{||\nabla_{z_{c^{+}}}\mathcal{L}||}{\sum_{c^{-}\neq c^{+}}^{C}||\nabla_{z_{c^{-}}}\mathcal{L}||}\approx\frac{1-\sigma(z_{c^{+}})}{(C-1)\cdot\mathbb{E}[\sigma(z_{c^{-}})]}\propto\frac{1}{C\cdot\epsilon}.$ (3)
@@
2) Hard Negative Gradient Dilution. The massive negative space leads to gradient dominance by easily classified negatives rather than informative hard negatives. Let $\mathcal{H}$ denote the set of hard negative classes with $\mathbb{E}[\sigma(z_{c^{h}})]=\epsilon^{h}$ for $c^{h}\in\mathcal{H}$. The ratio of hard negative gradients to total negative gradients is:
@@
Figure 3: Overview of the CQ-DINO framework for vast vocabulary object detection. Key components: (1) Learnable category queries enhanced with hierarchical tree construction for semantic relationship modeling; (2) Image-guided query selection that identifies the most relevant category queries; (3) Feature enhancer and cross-modality decoder (adapted from GroundingDINO [22]), processing object queries with contrastive alignment between object and selected category queries.
@@ 3.2 CQ-DINO
1) Learnable category queries with correlation encoding. We initialize learnable category queries $Q_{cat}\in\mathbb{R}^{B\times C\times D}$ using the OpenCLIP [26] text encoder, where $B$ is the batch size, $C$ denotes the total number of categories, and $D$ is the embedding dimension. These category queries enable flexible encoding of category correlations through self-attention or hierarchical tree construction (Sec. 3.4).
@@
3) Feature enhancer and cross-modality decoder. The selected category queries $Q_{cat}^{\prime}\in\mathbb{R}^{B\times C^{\prime}\times D}$ and image features $F_{img}$ are processed through GroundingDINO [22] components. First, the feature enhancer module fuses the category queries and image features. Next, object queries are generated through language-guided query selection. Finally, detection outputs are produced by the cross-modality decoder and contrastive alignment between object queries and selected category queries.
@@ 3.3 Image-Guided Query Selection
The core innovation of CQ-DINO is our image-guided category selection module, illustrated in Fig. 5. This module employs cross-attention between category queries $Q_{cat}$ and image features $F_{img}$. Here, $Q_{cat}$ serves as queries (Q), while $F_{img}$ provides keys (K) and values (V). The cross-attention layer establishes category-to-image correlations through similarity computation. Then, we apply TopK selection to retain only the top-$C^{\prime}$ categories ($C^{\prime}\ll C$) based on activation values. We supervise this selection using Asymmetric Loss [29], which serves as a multi-class classification loss.
@@
$\frac{\rho^{\prime}}{\rho}=\frac{\sum_{c^{-}}^{C}||\nabla_{z_{c^{-}}}\mathcal{L}||}{\sum_{c^{-}\in S}^{C^{\prime}}||\nabla_{z_{c^{-}}}\mathcal{L}||}\approx\frac{C}{C^{\prime}}.$ (5)
For a typical setting ($C>10^{4},~{}C^{\prime}=100$), this achieves a 100×100$\times$100 × gradient rebalancing factor. Our design provides three benefits: 1) Gradient rebalancing. By filtering out easy negative categories, our selection module improves the influence of gradients from positive examples. 2) Adaptive hard negative mining. The selection mechanism ensures that retained negative categories exhibit high semantic relevance to the image content, naturally implementing hard negative mining. 3) Scalable computation. Processing only $C^{\prime}$ categories reduces memory consumption and computational cost, making the framework scalable to extremely large vocabularies.

### v4

@@ Abstract
With the exponential growth of data, traditional object detection methods are increasingly struggling to handle vast vocabulary object detection tasks effectively. We analyze two key limitations of classification-based detectors: positive gradient dilution, where rare positive categories receive insufficient learning signals, and hard negative gradient dilution, where discriminative gradients are overwhelmed by numerous easy negatives. To address these challenges, we propose CQ-DINO, a category query-based object detection framework that reformulates classification as a contrastive task between object queries and learnable category queries. Our method introduces image-guided query selection, which reduces the negative space by adaptively retrieving the top-K relevant categories per image via cross-attention, thereby rebalancing gradient distributions and facilitating implicit hard example mining. Furthermore, CQ-DINO flexibly integrates explicit hierarchical category relationships in structured datasets (e.g., V3Det) or learns implicit category correlations via self-attention in generic datasets (e.g., COCO). Experiments demonstrate that CQ-DINO achieves superior performance on the challenging V3Det benchmark (surpassing previous methods by 2.1% AP) while maintaining competitiveness on COCO. Our work provides a scalable solution for real-world detection systems requiring wide category coverage.
33footnotetext: Corresponding author: Yongchao Xu <yongchao.xu@whu.edu.cn>44footnotetext: Source Code: https://github.com/FireRedTeam/CQ-DINO
1 Institute of Artificial Intelligence, School of Computer Science, Wuhan University
2 Xiaohongshu Inc.           3 Hubei Luojia Laboratory
11footnotetext: {zhichaosun, yongchao.xu}@whu.edu.cn22footnotetext: {huhuazhang, mayidong, tangshen}@xiaohongshu.com {liugang.spl, nemochen89, yaoohu}@gmail.com
@@ 1 Introduction
With the rapid expansion of data, developing a robust AI system capable of large-scale object detection has become essential. This necessity is driven by the increasing complexity and diversity of real-world applications, where AI must manage an extensive vocabulary and dynamic environments. Vast vocabularies inherently present hierarchical category structures, as illustrated by classification datasets like ImageNet deng2009imagenet and Bamboo zhang2022bamboo . Recent detection benchmarks, such as V3Det wang2023v3det , which feature 13,204 object categories organized in hierarchical structures, highlight the magnitude of this challenge. While object detection has witnessed significant advancements fasterrcnn ; glip ; gdino ; fu2025llmdet , scaling effectively to vast vocabularies remains a substantial challenge.
Category prediction mechanisms can be broadly categorized into four types, as illustrated in Fig. 1. Classification head-based methods employ feed-forward networks (FFNs) with sigmoid activation and Focal Loss ross2017focal optimization. These approaches perform well on benchmarks with limited categories such as COCO coco (80 categories) and Objects365 shao2019objects365 (365 categories), but face fundamental challenges and scalability issues in vast vocabulary settings. Text-prompted contrastive methods leverage Vision-Language Models (VLMs) to encode target categories as text inputs, achieving strong open-vocabulary detection ovcoco performance. However, in vast vocabulary scenarios, text input sequences become prohibitively long, necessitating the splitting of category lists across multiple inference passes. This substantially limits their practical scalability and efficiency. Language model-generated methods approach open-ended generateu object detection by using language models to generate category labels without predefined candidate sets. Although these methods theoretically enable the detection of arbitrary object categories, they generally lack mechanisms to control the granularity of generated labels, which can result in substantial misalignment with practical detection requirements.
@@
To tackle these challenges, we introduce Category Query-based DINO (CQ-DINO), a novel architecture that encodes categories as learnable query embeddings. Our approach centers on image-guided query selection, which identifies relevant categories via category-to-image similarity through cross-attention. The key insight driving our method is that dynamic sparse category selection significantly reduces the negative search space. This selection mechanism provides three crucial benefits: (1) balancing the ratio between positive and negative gradients, (2) performing implicit hard mining by selecting the most similar categories, and (3) reducing memory and computational costs, making the framework scalable to extremely large vocabularies. Selected category queries interact with image features to generate object queries. From these object queries, bounding boxes are predicted through a cross-modality decoder. The final classifications are obtained using contrastive alignment between the object queries and the category queries. Unlike traditional classification head-based methods, our category query representation offers greater flexibility by naturally encoding inter-category relationships. For structured datasets with explicit category hierarchies like V3Det wang2023v3det , we leverage the inherent tree structure to construct hierarchical category queries with an adaptive weighting mechanism that balances local and hierarchical features. For datasets without explicit hierarchies (e.g., COCO coco ), we employ self-attention mechanisms to implicitly learn the correlations between categories.
We evaluate CQ-DINO on both the vast vocabulary V3Det dataset wang2023v3det and the standard COCO benchmark coco . Our approach surpasses previous state-of-the-art methods on V3Det while maintaining competitive results on COCO compared to DETR-based detectors. Our method benefits from vast vocabulary detection while maintaining competitive results in limited vocabulary scenarios.
@@ 2.1 Vast Vocabulary Object Detection
The progression of object detection benchmarks reflects a steady growth in category vocabulary, evolving from relatively small vocabulary datasets such as PASCAL VOC everingham2015pascal (20 classes) and COCO coco (80 classes), to larger vocabulary benchmarks, including Objects365 shao2019objects365 (365 classes) and Open Images openimage (600 classes). Recently, Wang et al. wang2023v3det introduce V3Det, the first vast vocabulary object detection dataset, comprising 13,204 hierarchically structured categories. This unprecedented scale poses significant challenges in terms of scalability and representation.
Recent progress has been driven by the V3Det Challenge v3detchallenge , yielding several methodological advancements. MixPLv2 proposes a semi-supervised framework that combines labeled V3Det data with unlabeled Objects365 shao2019objects365 images through pseudo-labeling. RichSem-DINO-FocalNet enhances detection robustness by integrating the RichSem-DINO RichSem-DINO framework with a FocalNet-Huge backbone FocalNet pretrained on Objects365. Most recently, Prova chen2024comprehensive introduces multi-modal image-text prototypes specifically optimized for V3Det’s fine-grained classification. However, these methods still rely on fundamentally similar classification architectures, raising questions about their effectiveness and scalability for even larger vocabularies beyond tens of thousands of categories.
@@ 2.2 Object Detectors
Most object detection frameworks employ feedforward networks (FFNs) as fixed classification heads for category prediction. In two-stage detectors fasterrcnn ; cascadercnn ; he2017mask , region proposals are generated in a class-agnostic manner by Region Proposal Networks (RPNs), followed by FFN-based region classifiers. By contrast, one-stage detectors ross2017focal ; fcos ; duan2019centernet predict bounding boxes and categories in a single step. Transformer-based methods, such as DETR detr , reformulate detection as a set prediction problem using learnable queries. Although DETR achieves an elegant end-to-end design, it suffers from slow convergence. Subsequent works zhu2021deformable ; anchor ; dabdetr ; dndetr ; zhang2023dino ; pu2023rank ; liu2023detection ; hou2024relation mitigate these limitations. For example, Deformable DETR zhu2021deformable proposes multi-scale deformable attention for sparse spatial sampling. DINO zhang2023dino improves performance via contrastive query denoising and mixed query selection. Despite these architectural advances, current methods fundamentally rely on classification heads with activation functions, typically optimized with Focal Loss ross2017focal or Cross-Entropy Loss cross-entropy . This design inherently constrains scalability and presents optimization challenges when extending to vast vocabulary detection scenarios.
Text-prompted Contrastive Methods.
@@
Vision-language models have advanced the seamless integration of visual and textual modalities for open-vocabulary object detection. These methods encode target categories as text inputs and align visual and textual representations. For instance, GLIP glip pioneers the use of contrastive learning between image regions and textual phrases. Grounding DINO gdino further improves cross-modal alignment through early fusion of vision and textual features. Similarly, DetCLIP yao2022detclip and RegionCLIP zhong2022regionclip leverage image-text pairs with pseudo-labels to enhance region-level semantic understanding and improve generalization. Despite these advances, text-prompted methods face scalability bottlenecks due to the limited capacity of text token inputs during inference. For example, GLIP glip and Grounding DINO gdino restrict input prompts to approximately 128 tokens per pass, allowing for about 40 categories simultaneously. Thus, detecting all categories in a vast vocabulary benchmark like V3Det wang2023v3det (13,204 classes) would require over 331 sequential inference passes per image. This constraint renders current text-prompted methods computationally inefficient and impractical for real-time or large-scale detection.
Language Model-Generated Methods.
@@
Advances in multimodal large language models (MLLMs) have inspired detection methods leveraging their visual understanding and generative abilities. Some MLLMs internvl ; bai2023qwen ; gemini show preliminary object detection abilities but exhibit limited localization precision and recognition granularity. To mitigate these limitations, recent works generateu ; fu2025llmdet ; jiang2024chatrex ; wang2023v3det ; glee use LLMs primarily as category generators rather than direct detectors. For instance, GenerateU generateu employs a T5-based decoder t5 to generate category names from visual features, reframing detection as text generation. Similarly, LLMDet fu2025llmdet and ChatRex jiang2024chatrex utilize instruction-tuned LLMs to predict object categories from image features. While these generative approaches enable open-ended detection, they often produce inconsistencies due to limited controllability over label granularity. For example, given an image region of a “Persian cat”, the model may generate the generic term “cat”, causing semantic ambiguity and reduced accuracy for fine-grained detection tasks.
2.3 Category Query-based Methods
Learnable queries was popularized in computer vision by DETR detr , marking a paradigm shift from fixed architectural components to task-adaptive representations. Queries serve as learnable embeddings that interact with the visual feature space, enabling the model to capture complex, task-specific patterns. This design has since been adopted in diverse domains, including classification liu2021query2label , segmentation cheng2021per , and multimodal learning li2023blip . Among these developments, category queries represent an innovation introduced by Query2Label liu2021query2label . Rather than relying on fixed classification heads, Query2Label proposed learnable category embeddings to capture category-specific features. Subsequent works such as ML-Decoder ridnik2023ml have outperformed conventional classification methods. The effectiveness of category queries in classification motivated their extension to dense prediction tasks. For example, CQL xie2023category applies them to human–object interaction classification, ControlCap zhao2024controlcap uses them for semantic guidance in region captioning, and RankSeg he2022rankseg integrates them into semantic segmentation, dynamically selecting the top-$k$ most relevant classes during inference. This selective querying reduces the effective search space, improving computational efficiency and segmentation accuracy. While prior works have explored category queries in various contexts, our work addresses a fundamentally different challenge specific to vast vocabulary scenarios. We provide, to the best of our knowledge, the first systematic theoretical analysis of gradient dilution issues that arise when dealing with vast category vocabularies, which motivates our image-guided query selection design. Moreover, we introduce a hierarchical tree construction strategy that explicitly models category correlations, enabling effective reasoning over deep semantic hierarchies in vast vocabulary datasets.
@@ 3.1 Challenges in Vast Vocabulary Detection
Existing methods for vast vocabulary object detection with $C$ categories ($C>10^{4}$), particularly those employing sigmoid-based classifiers with Focal Loss ross2017focal , face fundamental optimization challenges. We systematically analyze these issues through a simplified formulation using the Cross-Entropy Loss with sigmoid activation, revealing two critical limitations:
1) Positive Gradient Dilution.
In vast vocabulary detection, the gradient signal for positive classes is overwhelmed by the aggregated negative gradients. Let $z_{c}$ denote the logit for class $c$ and $y_{c}\in\{0,1\}$ be its ground-truth label. The gradient of the Cross-Entropy Loss $\mathcal{L}$ with respect to $z_{c}$ is:
@@
[…]
@@
[…]
@@
2) Hard Negative Gradient Dilution.
The massive negative space leads to gradient dominance by easily classified negatives rather than informative hard negatives. Let $\mathcal{H}$ denote the set of hard negative classes with $\mathbb{E}[\sigma(z_{c^{h}})]=\epsilon^{h}$ for $c^{h}\in\mathcal{H}$. The ratio of hard negative gradients to total negative gradients is:
@@
Figure 3: Overview of the CQ-DINO framework for vast vocabulary object detection. Key components: (1) Learnable category queries enhanced with hierarchical tree construction for semantic relationship modeling; (2) Image-guided query selection that identifies the most relevant category queries; (3) Feature enhancer and cross-modality decoder (adapted from GroundingDINO gdino ), processing object queries with contrastive alignment between object and selected category queries.
@@ 3.2 CQ-DINO
1) Learnable category queries with correlation encoding. We initialize learnable category queries $Q_{cat}\in\mathbb{R}^{B\times C\times D}$ using the OpenCLIP clip text encoder, where $B$ is the batch size, $C$ denotes the total number of categories, and $D$ is the embedding dimension. These category queries enable flexible encoding of category correlations through self-attention or hierarchical tree construction (Sec. 3.4).
@@
3) Feature enhancer and cross-modality decoder. The selected category queries $Q_{cat}^{\prime}\in\mathbb{R}^{B\times C^{\prime}\times D}$ and image features $F_{img}$ are processed through GroundingDINO gdino components. First, the feature enhancer module fuses the category queries and image features. Next, object queries are generated through language-guided query selection. Finally, detection outputs are produced by the cross-modality decoder and contrastive alignment between object queries and selected category queries.
@@ 3.3 Image-Guided Query Selection
The core innovation of CQ-DINO is our image-guided category selection module, illustrated in Fig. 5. This module employs cross-attention between category queries $Q_{cat}$ and image features $F_{img}$. Here, $Q_{cat}$ serves as queries (Q), while $F_{img}$ provides keys (K) and values (V). The cross-attention layer establishes category-to-image correlations through similarity computation. Then, we apply TopK selection to retain only the top-$C^{\prime}$ categories ($C^{\prime}\ll C$) based on activation values. We supervise this selection using Asymmetric Loss ridnik2021asymmetric , which serves as a multi-class classification loss.
@@
[…]
For a typical setting ($C>10^{4},~C^{\prime}=100$), this achieves a $100\times$ gradient rebalancing factor. Our design provides three benefits: 1) Gradient rebalancing. By filtering out easy negative categories, our selection module improves the influence of gradients from positive examples. 2) Adaptive hard negative mining. The selection mechanism ensures that retained negative categories exhibit high semantic relevance to the image content, naturally implementing hard negative mining. 3) Scalable computation. Processing only $C^{\prime}$ categories reduces memory consumption and computational cost, making the framework scalable to extremely large vocabularies.